# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v1

@@ 5.4.2. Grok3
In the first case, wherein the default and few-shot responses perfectly align with those of participants, the model maintains a similar rationale between conditions and adjusts some responses based on the details in the provided scenario. For example, Grok explicitly reported that, while the sensitivity ratings are similar in both cases, it modified its response to one of the provided questions: “The main deviation is in data retention (Question 3), where I predict a shift from ‘delete immediately’ to ‘keep temporarily’ due to the utility of the data (reminders) and the lower concern level.” This explanation shows that Grok is capable of producing sound responses and reasoning with or without information about the user.
@@
We see a similar pattern in the second case, but neither the default nor the few-shot responses match the human responses perfectly. This happens because, while the user chooses the same sensitivity score for both the example and the targeting scenario, they also select fewer privacy-enhancing choices in response to the targeting scenario. In this case, while the Grok’s reasoning is sensible, it is difficult for the model to predict the participant’s specific preferences.
@@
In the third case, Grok’s default results align more closely with the participant’s responses than its few-shot results. When Grok provides default responses, it happens to select options that overall align well with the participant. In this case, Grok explains that privacy-enhancing options are ideal for the scenario, noting specific scenario details such as the involvement of a minor who “may not fully grasp the implications of data collection”. It also makes logical concessions, such as preferring that collected data be retained for a predefined period, rather than no data being retained, because the user requested private data recording that must be kept for some time to serve its purpose.
Once Grok receives additional information through the few-shot prompt, it is able to compare the participant’s responses in the additional example scenario, and then proceeds to over-correct when responding to the current scenario. While Grok previously provided a rationale entirely based on the details of the scenario, it instead focuses its responses on meeting the preferences of the participant. For example, Grok specifically notes that it is attempting to balance both the high POS scores of the participant and their low scenario sensitivity rating. These differences, while entirely valid, appear to be difficult for the Grok to mediate. As a result, Grok selects responses that are too relaxed, more consistent with a low sensitivity rating, and ultimately aligns poorly with the participant’s preferences.
@@
In the fourth scenario, Grok’s performance improves when provided with additional information (i.e., the few-shot prompt) compared to its default responses. By default, Grok again conveys that privacy-enhancing options are ideal, citing the involvement of a minor and the fact that the collected data, the visual of a report card, is irrelevant to the robot’s task. While these choices and rationale are reasonable, they do not align with the responses of the participant.
@@
When Grok is provided with additional information, it then focuses its choices and reasoning on the participant’s preferences. For example, Groks states that the participant’s “moderate score on privacy as a right … indicates they may not always demand strict privacy measures if the context seems low-risk”. In conjunction with the participant’s low sensitivity rating of the scenario and relaxed privacy preferences in a scenario Grok deemed more sensitive (i.e., the scenario involved medical data), Grok chose responses that were more relaxed overall. As a result, Grok’s responses were in greater alignment with the participant than they had been previously when selected agnostic of the participant.
5.4.3. Mistral
Overall, we find that Mistral does not provide clear justifications for its responses and generally provides rationales that are disconnected from the details of the scenarios themselves. We again provide examples showcasing this behavior.
Cases 2 & 4
Among these cases, it is difficult to evaluate the models’ responses themselves. The rationales Mistral provided were framed as “You chose ’a’ because…”, reflecting that it was attempting to justify responses it previously gave, rather than those it had just generated. This overall poor reasoning could also be due to the low temperature and Mistral’s small model size capability. When we observed an improvement through few-shot prompting, the logic Mistral provided did not support privacy awareness; rather it was difficult to understand how Mistral arrived at responses with greater alignment. It is possible that the model is able to capture nuances of privacy but has difficulty conveying them in its logical statements. Alternatively, it is possible that privacy nuances and the preferences of participants are not captured and, instead, the model performed well by chance. In both cases, Mistral does not refer to the details in the targeting scenario and at times shows logical flaws: “In this scenario, the person may not want the robot to notify them about what will happen to the data, as this would potentially increase their privacy concerns and make them feel like they have less control over their personal information.” This example shows that when Mistral does attempt to justify its responses, it is capable of making clear logical errors.
5.4.4. Patterns Across Models
@@
For both Grok and Mistral, we see frequent privacy-preserving responses under default prompting. When provided with details about participants, we see that both models tend to mirror the responses from the example scenario during the targeting scenario. We calculated the percentage of scenarios in which this outcome occurs; Grok provides the exact same responses (i.e., both categorical and binary) in 42% of scenarios. Considering only binary responses, the percentage rises to 89.89%. A similar result occurs in the case of Mistral, with identical responses in 49.33% of scenarios and 73.89% of scenarios among binary responses only.
@@
While we observe that Grok provides identical responses to many of the examples it is provided, it acknowledges that these decisions are the same. It frequently refers to the preferences a participant previously expressed and purposefully suggests a similar response. A full analysis of Grok’s logic across all scenarios, participants and few-shot responses would be required to truly assess whether Grok is providing privacy-aware answers, but that analysis falls outside the scope of this work. This case study, however, does suggest that Grok is utilizing the details that it has been provided to generate preferable responses to the particular user in question.
@@ 6. Discussion
In this work, we explored the relationship between privacy orientation and preferred robot behaviors, compared human-preferred robot behaviors to those provided by LLMs, and further evaluated various prompting strategies in terms of their ability to improve LLM performance. Regarding participant privacy orientation, we show that specific subscales of the POS are predictive of preferences for specific robot behaviors. Through our comparative analysis of human-preferred and LLM-generated robot responses, we show a low to moderate level of conformance (i.e., micro and macro F1 score). In our exploration of prompting strategies, we saw a substantial increase in LLM performance through few-shot prompting, with more modest or no performance gains among the other strategies. Through these results, we identify three broad implications of this work.
6.1. Private Mindset
The results of our user evaluation suggest that participants were highly privacy oriented (i.e., score high on average across all POS subscales) and preferred robots to behave in a privacy-preserving manner, even in minimally sensitive private scenarios. The difference between the POS scores of our participants and those of Baruh and Cemalcılar (2014) suggests a potential shift in privacy concerns since the original data collection in 2013. Given our sample was geographically based in the United States, we suspect that several major national events that have occurred since 2013 may have altered the privacy sentiments of our US-based participants. These include the government surveillance revelations of Edward Snowden in 2013 (Greenwald et al., 2013), concerns of user manipulation and election interference by Cambridge Analytica in 2016 (Detrow, 2018), the Equifax data breach in 2017 (Electronic Privacy Information Center (2017), EPIC), among others. Events of this magnitude could be responsible for an erosion in public trust of government and corporate agencies that collect large amounts of data and have significant responsibility for the management of that data. While we cannot draw any direct conclusions about these events, it is quite clear that the general U.S. population has concerns about privacy. Prior work by McClain et al. (2023) has shown that more people in 2023 say they are not knowledgeable about what happens with their data and have little understanding of the laws and regulations that protect their privacy than compared to 2019. As users are continually required to make concessions about their privacy in order to use various technologies, only to have their privacy further violated time and again, it would be unsurprising if such a shift has occurred. As the scope of our study cannot identify evidence of an increase in privacy concerns broadly, it is also possible that our sample population is either skewed or biased by elements of our survey (e.g., introducing robot involvement in the survey prior to completion of the POS).
@@ 6.2. LLM Privacy Awareness
Our results show that LLMs generally provide privacy-preserving responses, but they are not necessarily privacy-aware. We observe a trend where LLMs provide increasingly private responses when they are provided higher sensitivity ratings; however, they often fail to capture the nuance of why a given scenario is sensitive or the appropriate response to that scenario. From our case studies, Section 5, we see that even the best performing model can ‘overcorrect’ during few-shot prompting when the details of one scenario lead it to incorrectly infer stringent privacy preferences in another. We also find that larger models, particularly the closed-source ones, perform better than open-source models. From our case studies, Mistral-7B, a small model that could run on edge devices, may not be able to provide valid reasoning based on user inputs. Llama-Scout-17B, however, performs similarly to larger models, potentially indicating a bottom line of model size for designing.
6.3. Design Implications towards LLM-powered Robots
@@
Prior work has already implemented LLMs and VLMs in social robots directly (Benjdira et al., 2023) or fed them with a specific persona or rules (Abbo et al., 2025; Bezabih et al., 2024). In our study, however, we see that prompting LLMs with the addition of user privacy orientation or perceived sensitivity of scenarios only minimally improves performance, if not impairing or creating no change in performance. Conversely, using few-shot prompting substantially improves LLM performance on average. These results imply that an LLM-powered social robot can utilize naturalistic interactions with a user to inform future privacy-sensitive responses to that same user.
Bridging the Gap between Human and Robot
@@
While few-shot prompting may be able to improve model conformance to user preferences, it is worth revisiting the methods used in our few-shot prompting strategy. In our work, due to practical limitations and the necessity of standardized inputs in our study design, we utilized human responses for example scenarios. In live scenarios, however, a robot will not have access to standardized multiple-choice behavior preferences. Instead, the robot will need to be capable of understanding the ongoing scenarios. For example, VLMs may be used to allow a robot to recognize its environment, as many researchers have already shown (Wake et al., 2025; Zhong et al., 2024; Rahimi et al., 2025). By interpreting the visual and sound-to-text inputs, users may then be able to prompt the LLMs to capture and reason about scenarios similar to those introduced in this paper.
@@
Additionally, the robot must be able to signal its data collection to users. For instance, robots may utilize privacy-preserving behaviors by default (e.g., minimize data collection and retention) and display both sound and text indicators (e.g., “video recording in progress”) on the screen when needed. Future user studies are needed in order to design these indicators. Once a robot is able to recognize a scenario and respond to it, the robot then needs to be capable of discerning user responses or preferences in order to conform to users over time. While this outcome could be achieved through a user-driven calibration phase, only learning over time would allow the robot to understand the unique nuances of individual and family privacy preferences. For future work, we aim to implement such a system in existing social robots and conduct in-lab or in-home evaluations.
@@
As discussed in Section 5, closed-source models generally perform better than open-source ones. Closed-source models, however, run over cloud API and, while most companies such as OpenAI do not use API user data for training by default (OpenAI, 2025), these data are stored temperately on cloud servers for debugging purposes. Given that prior works (Afroogh et al., 2024; Lockey et al., 2024; Chen et al., 2025) show that users have low levels of trust toward AI models’ data processing practices, using non-local models on social robots might cause privacy concerns. The goal of designing LLM-powered robots should account for this type of concern, and we, therefore, propose two potential solutions. First, if non-local model API calls are used, users should be clearly informed where their data is going and how their data will be used. Second, given the clear benefits of local models for user privacy (i.e., reduced data privacy concerns), future studies can be used to fine-tune these models, or distill them from open source models, such that they are sufficiently privacy aware to also mitigate interpersonal privacy concerns.
@@ 6.4. Limitations
There are two primary limitations presented in this work. First, participant POS scores do not align with prior work, which suggests a possible skew or bias in our sample (see Section 7.3). While many other scales exist that center on privacy concerns related to technology (see Section 2.1), many of these focus on either online privacy, privacy concerns of specific populations, or privacy broadly. We have not identified, however, scales that are specific to in-context human-robot interaction. Development of such a scale may be worth exploring further as future work. Secondly, while we have evaluated the performance of several models, identified strengths and weaknesses of those models, and explored strategies to improve model performance through prompting, we still cannot definitively conclude whether LLMs are truly privacy-aware. To do so, further analysis would be required involving additional qualitative analysis over a set of state-of-the-art models. For example, if we were able to analyze the logic of all models across a similar broad set of scenarios, we may be able to better identify the unique cases in which models align well with human preferences or fail to capture these norms, and what information is critical for model improvement.
@@ 7. Conclusion
In this work, we have evaluated the privacy awareness of ten state-of-the-art LLMs in the context of household human-robot interaction. While our results show that LLMs possess broad limitations in terms of their ability to conform to user privacy preferences, the prompting strategies we have presented may be able to bridge this gap. While the considerations of user privacy are varied and complex, the results of this work suggest that LLMs have the potential to act as rational privacy controllers for social robots capable of enhancing user privacy.
Appendix
7.1. Questions
@@ 7.2. Codebook
Only share or use data data when given permission 5
@@
7.3. POS Statistics Comparing to Prior work

### v2

@@ 5.4.2. Results
In the first case, wherein the default and few-shot responses perfectly align with those of participants, the model maintains a similar rationale between conditions and adjusts some responses based on the details in the provided scenario. For example, Grok explicitly reported that, while sensitivity ratings are similar in both cases, it modified its response to one of the provided questions: “The main deviation is in data retention (Question 3), where I predict a shift from ‘delete immediately’ to ‘keep temporarily’ due to the utility of the data (reminders) and the lower concern level.” This explanation shows that Grok is capable of producing sound responses and reasoning with or without information about the user.
@@
We see a similar pattern in the second case, but neither the default nor the few-shot responses match the human responses perfectly. This happens because, while the user chooses the same sensitivity score for both the example and the targeting scenario, they also select fewer privacy-enhancing choices in response to the targeting scenario. In this case, while Grok’s reasoning is sensible, it has difficulty in predicting the participant’s specific preferences.
@@
In the third case, Grok’s default results align more closely with the participant’s responses than its few-shot results. When Grok provides default responses, it happens to select options that overall align well with the participant. In one example, Grok explains that privacy-enhancing options are ideal for the scenario, noting specific scenario details such as the involvement of a minor who “may not fully grasp the implications of data collection”. Grok also makes logical concessions such as preferring that collected data be retained for a predefined period, rather than no data being retained, because the user in the scenario requested private data recording that must be kept for some time to serve its purpose.
Once Grok receives additional information through the few-shot prompt, it is able to compare the participant’s responses in the additional example scenario, and then proceeds to over-correct when responding to the current scenario. While Grok previously provided a rationale entirely based on the scenario’s details, it instead focuses its responses on meeting the participant’s preferences. For example, Grok specifically noted that it was attempting to balance both the high POS scores of the participant and their low scenario sensitivity rating. These differences, while entirely valid, appear to be difficult for Grok to mediate. As a result, Grok selected responses that were too relaxed, more consistent with a low sensitivity rating, and ultimately aligned poorly with the participant’s preferences.
@@
In the fourth scenario, Grok’s performance improves when provided with additional information (i.e., the few-shot prompt) compared to its default responses. By default, Grok again conveys that privacy-enhancing options are ideal, citing details like the involvement of a minor or the fact that the collected data is irrelevant to the robot’s task. While these choices and rationales are reasonable, they do not align with the responses of the participant.
@@
When Grok is provided with additional information, it then focuses its choices and reasoning on the participant’s preferences. For example, in one scenario Grok stated that a participant’s “moderate score on privacy as a right … indicate[d] they may not always demand strict privacy measures if the context seems low-risk.” In conjunction with the participant’s low sensitivity rating of the scenario and relaxed privacy preferences in a scenario Grok deemed more sensitive (i.e., the scenario involved medical data), Grok chose responses that were more relaxed overall. As a result, Grok’s responses were in greater alignment with the participant than they had been previously when selected agnostic of the participant.
5.4.3. Patterns Across Scenarios
@@
Among Grok’s rationales, we see frequent privacy-preserving responses under default prompting. When provided with details about participants, we see that Grok tends to mirror the responses from the example scenario during the targeting scenario. We calculated the percentage of scenarios in which this outcome occurs; Grok provides the exact same responses (i.e., both categorical and binary) in 42% of scenarios. Among only binary responses, the percentage rises to 89.89%.
@@
While we observe that Grok provides identical responses to many of the examples it is provided, it acknowledges that these decisions are the same. It frequently refers to the preferences a participant previously expressed and purposefully suggests a similar response. A full analysis of Grok’s logic across all scenarios, participants, and few-shot responses would be required to truly assess whether Grok is providing privacy-aware answers, but that analysis falls outside the scope of this work. These case studies, however, do suggest that Grok is utilizing provided details to generate appropriate and preferable responses to the particular user in question.
@@ 6. Discussion
In this work, we explored the relationship between privacy orientation and preferred robot behaviors, compared human-preferred robot behaviors to those provided by LLMs, and further evaluated various prompting strategies in terms of their ability to improve LLM performance. Regarding privacy orientation, we show that specific subscales of the POS are predictive of preferences for specific robot behaviors. Through our comparative analysis of human-preferred and LLM-generated robot responses, we show a low to moderate level of conformance (i.e., Micro and Macro F1 scores). In our exploration of prompting strategies, we saw a substantial increase in LLM performance through few-shot prompting, with more modest or no performance gains among other strategies. Through these results, we identify three broad implications of this work.
6.1. Privacy Mindset
The results of our user evaluation suggest that participants were highly privacy oriented (i.e., on average across all POS subscales) and preferred robots to behave in a privacy-preserving manner, even in minimally sensitive private scenarios. The difference between the POS scores of our participants and those of Baruh and Cemalcılar (2014) suggests a potential shift in privacy concerns since the original data collection in 2013. Given our sample was geographically based in the United States, we suspect that several major national events that have occurred since 2013 may have altered the privacy sentiments of our US-based participants. These include the government surveillance revelations of Edward Snowden in 2013 (Greenwald et al., 2013), concerns of user manipulation and election interference by Cambridge Analytica in 2016 (Detrow, 2018), and the Equifax data breach in 2017 (Electronic Privacy Information Center (2017), EPIC), among others. Events of this magnitude could be responsible for an erosion in public trust of government and corporate agencies that collect large amounts of data and have significant responsibility for the management of that data. While we cannot draw any direct conclusions about these events, it is quite clear that the general U.S. population has concerns about privacy. Prior work by McClain et al. (2023) has shown that more people in 2023 say they are not knowledgeable about what happens with their data and have little understanding of the laws and regulations that protect their privacy than compared to 2019. As users are continually required to make concessions about their privacy in order to use various technologies, it would be unsurprising if such a shift has occurred. As the scope of our study cannot identify evidence of an increase in privacy concerns broadly, it is also possible that our sample population is either skewed or biased by elements of our survey (e.g., introducing robot involvement in the survey prior to completion of the POS).
@@ 6.2. LLM Privacy Awareness
Our results show that LLMs generally provide privacy-preserving responses, but they are not necessarily privacy-aware. We observe a trend where LLMs provide increasingly private responses when they are provided higher sensitivity ratings; however, they often fail to capture the nuance of why a given scenario is sensitive or the appropriate response to that scenario. From our case studies (see Section 5), we see that even the best-performing model can ‘overcorrect’ during few-shot prompting when the details of one scenario lead the model to incorrectly infer stringent or weak privacy preferences in another. We also find that larger models, particularly the closed-source ones, perform better than open-source models. A cursory examination of the rationale of a small model that can be run on edge devices (i.e., Mistral), suggests that such models may not be able to provide valid reasoning based on user inputs. Llama 4 Scout, however, performs similarly to larger models, potentially indicating a bottom line of model size for designing.
6.3. LLM-Powered Robot Design Implications
@@
Prior works have already implemented LLMs and VLMs in robots directly (Benjdira et al., 2023) or provided them with specific personas or rules (Abbo et al., 2025; Bezabih et al., 2024). In our study, however, we see that prompting LLMs with the addition of user privacy orientation or perceived sensitivity of scenarios only minimally improves performance, if not impairing performance or creating no change. Conversely, using few-shot prompting substantially improves LLM performance on average. These results imply that an LLM-powered robot can utilize naturalistic interactions with a given user to inform future privacy-sensitive responses to that same user.
Practical Considerations of LLM Decision Making
@@
While few-shot prompting may be able to improve model conformance to user preferences, it is worth revisiting the methods used in our few-shot prompting strategy. In our work, due to practical limitations and the necessity of standardized inputs in our study design, we utilized a set of curated behaviors as responses to our example scenarios. In live scenarios, however, a robot will not have access to standardized multiple-choice behavior preferences. Instead, the robot will need to be capable of understanding ongoing scenarios and the unique responses those scenarios may require. For example, VLMs may be used to allow a robot to recognize its environment, as many researchers have already shown (Wake et al., 2025; Zhong et al., 2024; Rahimi et al., 2025). A VLM-equipped robot can subsequently reason about its environment, including details relevant to privacy, and appropriately tailor its response to the context of a scenario.
@@
Additionally, the robot must be able to signal its data collection to users. For instance, robots may utilize privacy-preserving behaviors by default (e.g., minimize data collection and retention) and display both sound and text indicators (e.g., “video recording in progress”) on the screen when needed. Future user studies are needed in order to design these indicators. Once a robot is able to recognize a scenario and respond to it, the robot then needs to be capable of discerning user responses or preferences in order to conform to users over time. While this outcome could be achieved through a user-driven calibration phase, only learning over time would allow the robot to understand the unique nuances of individual and family privacy preferences and how such preferences may change over time. For future work, we aim to implement such a system in an existing robot platform and conduct in-lab or in-home evaluations.
@@
As discussed in Section 5, closed-source models generally perform better than open-source ones. Closed-source models, however, run over cloud API and, while most companies such as OpenAI do not use API user data for training by default (OpenAI, 2025), these data are stored temperately on cloud servers for debugging purposes. Given that prior works (Afroogh et al., 2024; Lockey et al., 2024; Chen et al., 2025) show that users have low levels of trust toward AI models’ data processing practices, using non-local models on robots might cause privacy concerns. The goal of designing LLM-powered robots should account for this type of concern, and we, therefore, propose two potential solutions. First, if non-local model API calls are used, users should be clearly informed where their data is going and how their data will be used. Second, given the clear benefits of local models for user privacy (i.e., reduced data privacy concerns), future studies can be used to fine-tune these models, or distill them from open source ones, such that they are sufficiently privacy aware to also mitigate interpersonal privacy concerns.
@@ 6.4. Limitations
There are two primary limitations presented in this work. First, participant POS scores were higher on average in our study than in prior work. It is not clear whether this result reflects a true population change in privacy orientation over the last decade, as described earlier, or a skew or bias in our sample (see Section C). While many other scales exist that center on privacy concerns related to technology (see Section 2.1), many of these focus on either online privacy, concerns of specific populations, or privacy broadly. We have not identified, however, scales that are specific to in-context human-robot interaction. Development of such a scale may be worth exploring further as future work. Secondly, while we have evaluated the performance of several models, identified strengths and weaknesses of those models, and explored strategies to improve model performance through prompting, we still cannot definitively conclude whether LLMs are sufficiently privacy-aware. To do so, further qualitative analysis of state-of-the-art models is required. For example, if we are able to analyze the logic of all models across our scenarios, we may be able to better identify the unique cases in which models succeed or fail to meet user privacy preferences. In doing so, we can then gain a better understanding of whether LLMs are truly capable of privacy-aware decision making.
@@ 7. Conclusion
In this work, we have evaluated the privacy awareness of ten state-of-the-art LLMs in the context of household human-robot interaction. Our results show that LLMs possess broad limitations in terms of their ability to conform to user privacy preferences, but the prompting strategies we have presented may be able to bridge this gap. While the considerations of user privacy are varied and complex, the results of this work suggest that LLMs have the potential to act as rational privacy controllers for social robots capable of enhancing user privacy.
Appendix A Questions
@@ Appendix B Codebook
Only share or use data when given permission 5
@@
Appendix C POS Statistics Comparing to Prior work